AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  DESPEC  ELOG logo
Entry  Wed Oct 15 09:41:24 2025, JB, GB, MP, 207Bi source threshold tests IMG_5796.jpgIMG_5797.jpgIMG_5798.jpg
    Reply  Thu Oct 23 12:22:31 2025, JB, GB, MP, AM, NH, 207Bi source threshold tests 8x
       Reply  Fri Oct 24 10:31:59 2025, JB, GB, MP, AM, NH, 207Bi source threshold tests 6x
          Reply  Fri Oct 24 12:29:02 2025, JB, GB, MP, AM, NH, 207Bi source threshold tests 11x
Message ID: 723     Entry time: Fri Oct 24 10:31:59 2025     In reply to: 721     Reply to this: 724
Author: JB, GB, MP, AM, NH 
Subject: 207Bi source threshold tests 
> > 
> > We have dismounted the AIDA stack and bPlast and realigned the cables so that they can fit into the adapters.
> > 
> > The plan for the next couple days is the following:
> > 
> > - Alignment -- done
> > - Remount cable -- done
> > - Fix position w/o bPlast mounted -- done
> > - Remount bracket for the last stage -- done
> > - Remount AIDA 1 & 2 and the 207Bi source with the source holder -- done
> > - Mount the snout and take data -- we wait till next week Monday when the heat exchanger is active again. (ongoing)
> > 
> > The current setup at the end of work is that the snout is mount with a bPlast (uncabled) and 2 AIDA layers at the regular spaced position with a 207Bi source down stream, as can be seen in the previous elog. The source is ~1 cm away from the middle AIDA wafer.
> > 
> > Currently what is cabled:
> > 
> > - Signal out from the detector
> > - Bias daisy chain
> > 
> > Still to be cabled:
> > 
> > - HV cable
> > - Grounding
> > - Pulser
> > 
> > Additional info on the source position:
> > - the source is at ~9.6 mm distance from the downstream detector
> > - the source is placed at the centre of the middle wafer: 12.25 cm in x, 4.75 cm in y
> > - the radius of the source is ~1.5 mm
> 
> The Messhuette has been cleaned and equipment has been rearranged. If there are any broken or missing components be sure to follow up with the FRS group (as I believe they have been cleaning here).
> 
> We are resuming the work on the source measurements as the cooling water is back. Some of the valves still look closed so we are investigating if the water is flowing to the setup.
> 
> 14:31 Temperature on start up OK - attachment 2.
> 
> 15:41 We reset the ASIC control registers to a setup used in Feb 2024 (2025Feb24-17.14.52) to match the full AIDA setup (it was using BB7 initially in aida10).
> 
> The old bias scheme is obsolete and the bias is now done via the Ohmic side.
> 
> 16:47 We have look at some initial spectra with the 207Bi 3-8 shows the ADC channels and waveforms of the n+n and p+n sides. Bias and grounding developed by Nabiel have been done.
> 
> We are going to wait for the system to stabilise over night so that we start taking data tomorrow. Then we can analyse the results and see what we get :) Some channels are clearly broken which can be seen in the waveforms eg. aida16.

11:31 The detector was turned off last night to put it into a safe state. We return to make some comparison of the noise w.r.t the tests performed by Nabil and co. 

email snippet:

"
I would suggest that you first quantify current noise performance (slow comparators -> 0x64 to reduce rate, measure pulser peak width) for each FEE64 (use layouts 3 and 4) and compare to Nabil's most recent test results. Looking at last night's 1.8.W spectra I would assume that they would be significantly worse but best to check.

I would then suggest you carefully check

    Grounding configuration (both internal and external to the snout) used Nabil's most recent tests has actually been re-established
    Grounding of the bPlas cabling
"

We are going to restart the DAQ and set the LEC thresholds to 0x64 and compare the peak FWHM with a pulser.

11:50 We started up the DAQ temps and bias look OK.

12:00 We set the thresholds on each ASIC in the slow comparator LEC to 0x64. This should correspond to a high threshold. For some reason the y-strips are still noisy (wrong polarity in the register ?)
Attachment 1: Screenshot_from_2025-10-24_11-56-06.png  204 kB  Uploaded Fri Oct 24 12:00:56 2025  | Hide | Hide all
Screenshot_from_2025-10-24_11-56-06.png
Attachment 2: Screenshot_from_2025-10-24_12-00-30.png  156 kB  Uploaded Fri Oct 24 12:00:56 2025  | Hide | Hide all
Screenshot_from_2025-10-24_12-00-30.png
Attachment 3: Screenshot_from_2025-10-24_11-59-35.png  156 kB  Uploaded Fri Oct 24 12:00:56 2025  | Hide | Hide all
Screenshot_from_2025-10-24_11-59-35.png
Attachment 4: Screenshot_from_2025-10-24_11-55-02.png  125 kB  Uploaded Fri Oct 24 12:00:56 2025  | Hide | Hide all
Screenshot_from_2025-10-24_11-55-02.png
Attachment 5: Screenshot_from_2025-10-24_11-54-09.png  40 kB  Uploaded Fri Oct 24 12:00:56 2025  | Hide | Hide all
Screenshot_from_2025-10-24_11-54-09.png
Attachment 6: Screenshot_from_2025-10-24_11-53-59.png  160 kB  Uploaded Fri Oct 24 12:00:56 2025  | Hide | Hide all
Screenshot_from_2025-10-24_11-53-59.png
ELOG V3.1.3-7933898